OpenStudy (anonymous):

Wat is the equation of a sphere? Does it indicate that a point is on the surface of the sphere, or that a point is in the sphere itself?

jimthompson5910 (jim_thompson5910):

Point that is on the surface of the sphere.

OpenStudy (anonymous):

Depends whether u are a physicist or a math person..:-)

OpenStudy (anonymous):

I am talkimg about math

OpenStudy (anonymous):

i dont get this

OpenStudy (anonymous):

Says the champion :)

OpenStudy (anonymous):

its on the surface, if it were inside the sphere the equation would be like\[(x-a)^2+(y-b)^2+(z-c)^2<r^2\]

OpenStudy (anonymous):

That is not an equation.

jimthompson5910 (jim_thompson5910):

but that's not an equation anymore dodomaroc, it's an inequality.

OpenStudy (anonymous):

But otherwise you are correct

OpenStudy (anonymous):

Do you mean a formula for its volume or its surface area?

OpenStudy (anonymous):

Neither.

jimthompson5910 (jim_thompson5910):

neither volume nor surface area, this equation simply shows you which points lie on the surface of the sphere and when graphed, the visual is a sphere.

myininaya (myininaya):

do you how we say (x-h)^2+(y-k)^2=r^2 is the equation of a circle right?

OpenStudy (anonymous):

Ya

OpenStudy (anonymous):

n-sphere..

OpenStudy (anonymous):

wats an n-sphere?

myininaya (myininaya):

i bet we could use the distance formula to show the equation of a sphere like we can use the distance formula to show the equation of a circle

myininaya (myininaya):

but we have 3 dimensions and instead of 2

myininaya (myininaya):

so distance \[d=\sqrt{(x-x_1)^2+(y-y_1)^2+(z-z1)^2}\]

OpenStudy (anonymous):

hmmm. interesting

myininaya (myininaya):

dang it stupid subscripts

OpenStudy (anonymous):

So does this shows the distance between what???

myininaya (myininaya):

\[d=\sqrt{(x-x_1)^2+(y-y_1)^2+(z-z_1)^2}\]

myininaya (myininaya):

\[d^2=(x-x_1)^2+(y-y1)^2+(z-z1)^2\] center is \[(x_1,y_1,z_1)\] its the distance from a point on the sphere to the center of the sphere

OpenStudy (anonymous):

r^2 = sum i=1 to n+1 (x_i-c_i)^2